### v2

# Light-induced ultrafast magnetization dynamics in van der Waals antiferromagnetic CrSBr

Abstract
The magnetization dynamics driven by the femtosecond laser pulse of antiferromagnet van der Waals semiconductor CrSBr is studied within time-dependent density functional theory. We investigate the effect of laser fluence as well as the excitation frequency on the ultrafast dynamics of spins. In low fluence, the local magnetic moment of Cr increases when the laser frequency is below the band gap, whereas it decreases when the laser frequency is above the band gap. In high fluence, we find strong demagnetization independent of excitation frequency. We find that the ultrafast demagnetization in CrSBr is dominated by intralayer and interlayer optical intersite spin transfer, and spin flip via the spin-orbit coupling plays a minor role. Our results reveal hole excitation in the low-fluence regime and electron excitation in the high-fluence regime. Further, we investigate the effect of the external static magnetic field on the dynamics. We demonstrate even and odd high harmonic generation in the local magnetic moments and electric current, respectively. The external magnetic field results in an out-of-plane charge current via the inverse spin Hall effect as well as odd harmonics in the local magnetic moment dynamics as a result of breaking the time-reversal symmetry.
Introduction.— Van der Waals (vdW) magnets provide a promising platform for applications in both quantum and classic information processing [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. They exhibit atomically sharp interfaces enabling heterostructures engineering [11] as well as exceptional tunability through electric [12, 13, 14], magnetic [15, 16], mechanical [17, 18, 19, 18], and optical means [17, 9]. Particularly, vdW magnets can exhibit semiconducting electronic structure, facilitating strong couplings between magnetic structures and optical excitations [20, 15, 16, 21]. CrSBr is a prototype antiferromagnetic semiconductor showing strong coupling between microwave-range magnon and optical-range excitons [15, 16, 22], which provides an ideal platform for bridging microwave and optical regimes [23, 24, 25], positioning as a new frontier for spintronics, magnonics, and quantum information technologies.
Understanding and controlling vdW magnetic materials are crucial [26, 27, 28]. Ultrafast laser pulses have become a crucial and powerful means since the observation of the pulse-induced demagnetization in nickel within a sub-picosecond time scale [29], providing an ultrafast and non-thermal way to probe and control magnetic materials [28]. It has been shown that, when CrSBr is excited by ultrafast laser pulse, intriguing phenomena like periodically modulated exciton energy, nonlinear magnon frequency mixing, fast spin wavepacket transport, and even extreme high harmonic generations [15, 16, 30, 31, 32, 33], have been observed.
However, a microscopic understanding of how coherent spin dynamics is excited is still missing, especially when the nonequilibrium optical excitations are concerned, where magnetization from the excitations and local -orbital electrons are entangled. In this work, we present a full ab initio description of the ultrafast magnetization dynamics in CrSBr using real-time TDDFT (rt-TDDFT). Indeed, TDDFT gives a microscopic understanding of the ultrafast demagnetization, when the magnetization amplitude changes and the phenomenological Landau-Lifshitz-Gilbert formalism of magnetization with fixed amplitude, is not valid anymore. Here, we use TDDFT to systematically study the effect of the laser fluence, frequency, and magnetic fields on the magnetization dynamics of the local atomic moments in CrSBr. While the previous experimental and theoretical studies always showed demagnetization of the local moments in magnetic metals, here we observe an increase in these moments when the system is excited with a low fluence laser pulse with below gap central frequency. With low fluence laser, we find that the local moments show strong dependence on the laser central frequency. At high fluence, we find demagnetization of the local magnetic moments that is independent of the laser frequency.
Formalism.— CrSBr is an A-type vdW antiferromagnetic semiconductor where Cr atoms have nonzero local magnetic moment. The intralayer and interlayer magnetic orders are ferromagnetic and antiferromagnetic, respectively, with triaxial magnetic anisotropy and the easy axis being in the plane, Fig. 1 (a) [34]. It is air stable material with bulk Néel temperature K [22].
We employ the start-of-the-art rt-TDDFT method to study the ultrafast magnetization dynamics in bulk CrSBr excited by a laser pulse. The effect of laser pulse is captured by the time-dependent Kohn-Sham equation for spinor Kohn-Sham orbitals (using )
where is the vector potential of the laser pulse; is the speed of light; is the effective time-dependent Kohn-Sham potential, contributed by external potential from nuclei , Hartree potential , and exchange potential ; is the time-dependent Kohn-Sham magnetic field with being the external magnetic field and the exchange magnetic field; are Pauli matrices; and the last term describes spin-orbit coupling.
We employ adiabatic local spin density approximation (ALSDA) for exchange functional within full-potential augmented plane-wave method as implemented in the ELK code [35, 36]. The calculation is done in two stages; first, the ground state is obtained from ELK using static DFT calculations with local density approximation (LDA) exchange functional. We use DFT+U with eV, eV, and fully localized limit double counting scheme is applied. The ground-state band gap is eV and the magnetic moment for Cr atoms in agreement with literature [37, 38]. Next, the time-dependent calculation is performed using the ground-state Hamiltonian of the first step with a Gaussian femto-second laser pulse. In this study, the laser pulse is polarized along -direction traveling along -direction with wavelength of 1240 nm (below the band gap (), ) and 620 nm (above the band gap, ), full width at half maximum of 12 fs, and different laser pulse energies. The simulation time is 150 fs and a time step of 0.6 attosecond is chosen. The grid of vectors is chosen as . Since the wavelength of applied laser light is much larger than the supercell, we assume dipole approximation and disregard spatial dependence of the vector potential . In calculations with the external magnetic field, a static external magnetic field () is added to the Hamiltonian in both ground-state and time-dependent simulations.
These Kohn-Sham orbitals give rise to particle density and spin magnetization , so that the magnetic moment per unit cell is given by , integrated over the unit cell. Microscopic magnetization dynamics from purely electronic mechanisms is given by [39]
where is the total potential within the crystal, is the charge current along the -direction, and is the spin current along the -direction and -spin polarization, . Equation (2) demonstrates that magnetization dynamics is driven not only by external fields, but also by spin currents from optically excited carriers as well as spin-orbit coupling that is manifested as spatial derivatives of the potential. In this study, the ions are assumed fixed in the time-dependent calculations and we only focus on about 100 fs after the optical excitation where the most relevant dynamics is driven by the electronic system and this approximation is to a good extend valid.
Low fluence excitation.— Here we study the effect of the low-fluence () light on the magnetization dynamics. We find that the magnetization change shows a strong dependence on the light frequency. When the light frequency is below the band gap, the local magnetic moments of the Cr atoms are enhanced as shown in Fig. 1(a). This figure displays the time evolution of the magnetic moment of the Cr atom () in the layer that the magnetization points along the direction (which lies along the lattice vector ). The magnetic moments of Cr atoms in the other layer are opposite leading to zero net magnetization, guaranteed by the two-fold rotation symmetry about -axis. The laser pulse is illustrated by the black curve that lasts for about 12 fs. After the pulse, increases by about 2%. The blue solid line and red dotted line show the calculations with and without SOC, respectively. One can find that the effect of SOC is weak and does not change the qualitative behavior of the evolution of .
The increase of the magnetic moment arises from spin transfer between magnetic and nonmagnetic atoms. As shown in Fig. 1(b) calculated with SOC, the charge density of majority spin in Cr atom increases while that in S and Br atoms in the same layer decreases. For spin minority in Fig. 1(c), the trend is opposite. This indicates net spin transfer from S and Br to Cr, increasing the magnetic moment of the Cr atoms. In addition, we note that the reduction of spin in intralayer S and Br atoms is smaller than the increase of spin in Cr atom, suggesting that there is interlayer spin transfer. The spin transfer process is analogous to that in metallic materials where both magnetic and nonmagnetic atoms are present. However, in those systems, the spin transfer, termed optical inter-site spin transfer (OISTR) [40], typically induces demagnetization. Here we show that in the antiferromagnetic semiconductor, the OISTR can also increase the magnetic moment of magnetic atoms.
The spin transfer shows strong dependence on the light frequency. When the light frequency is above the band gap, the local magnetic moments of the Cr atoms are reduced or so-called demagnetized as shown in Fig. 1(e). Similar to the excitation with the below-gap frequency, the effect of SOC is weak, which slightly enhances demagnetization. The mechanism of the change of magnetic moment in Cr atom is similar. Figures 1(f) and 1(g) show that the averaged atomic spin density of S and Br decreases, whereas the spin density of Cr first increases and then reduces to amounts less than the equilibrium in the majority spin channel. In addition, the intralayer spin increase of Br and S is much weaker than the spin decrease in Cr atom, indicating that the interlayer spin transfer dominants the demagnetization.
High fluence excitation.— In the high fluence case (), we find that the qualitative behavior of spin dynamics shows weak dependence on the laser frequency. We focus on the case with below-gap frequency. With collinear antiferromagnetic order, the laser pulse does not induce a net magnetization. The magnetizations in neighboring layers are still opposite. Figure 2(a) plots the magnetic moment of the Cr atoms in one sublattice, where we find strong demagnetization on this sublattice in the presence of a laser pulse. One can see that the dynamics of starts with a slight enhancement, and then after a few fs, the demagnetization initiates. We observe about 50% demagnetization, in contrast to the low fluence cases, which shows weak dependence on the SOC. The effects of SOC start to manifest themselves near the end of the laser pulse, which introduces a weak correction that does not change the qualitative behavior.
The demagnetization in Cr atom leads to a reduction in the Néel vector despite a slight increase in the total magnetic moment of S and Br as shown in Figs. 2(b) and 2(c). In the absence of SOC, the spin flip scattering is forbidden. Therefore, the presence of demagnetization in one layer indicates the presence of interlayer spin transfer. In the presence of SOC, the demagnetization is slightly enhanced, indicating that the SOC-induced spin flip scattering can enhance the demagnetization.
Our further analysis of the time-dependent atomic spin densities are shown in Fig. 2 (b), (c), (f), and (g). Form Fig. 2 (b) and (f) one can see that the symmetric phase of the dynamics actually enhances the magnetic moments of the Cr atoms and this is the reason of the slight enhancement of the that was already discussed. However, the interlayer OISTR very soon dominates the dynamics resulting in fast demagnetization of . On the other hand, Fig. 2 (c) and (g) show a more lasting symmetric phase and therefore a stronger intralayer OISTR which results in demagnetization. Later on, the interlayer OISTR participates which is evident from the the nonsymmetrical phase of the dynamics. Here again, one may conclude that the interlayer and intralayer OISTR mechanisms are highly dependent on the spin channel and the excitation frequency. Spin-flip scatterings via SOC plays a small portion of the demagnetization process and comes into play with a delay.
Free carrier density.—Fig. 3 shows the densities of the excited free carriers i.e. the charge density in the interstitial region for both high and low fluence excitations with AG and BG frequencies. As can be seen in Fig. 3(a) in the low fluence excitation, the excitation results in hole generation while in the high fluence excitation, the excitation results in electron generation after the laser pulse is gone as can be seen in Fig. 3(b). This does not depend on the excitation frequency which is rather surprising in the low fluence case regarding the strong frequency dependence of the atomic spin densities and local magnetic moment of Cr atoms. In both figures, one can see a much larger amplitude of the variation in the case of BG excitation. This is because the AG and BG excitations are dominated by resonant and non-resonant transitions, respectively [41].
Effect of external magnetic field.— In the presence of the external magnetic field, the dynamics imply the generation of transverse magnons. The dynamics of the magnetic moment of Cr atoms along the x and z directions and in the presence of an external magnetic field are shown in Fig. 1 (d) and (h) for the AG and BG excitations, respectively. We note that the dynamics along the y direction in the presence of the external magnetic field is very similar to the dynamics in the absence of the external magnetic field shown in Fig. 1 (a) and therefore are not shown for brevity. We observe the initiation of low frequency oscillations of the magnetic moments along x and z direction in the presence of the external magnetic field indicating magnon generation. In addition, while the dynamics are similar for AG and BG excitations in the absence of the external magnetic field, they are quite different in the presence of the external magnetic field. The same conclusion can be obtained for the high fluence excitation, i.e. external magnetic field causes transverse magnon generation when the system is optically excited.
Electron dynamics and HHG.— In Fig. 4, we plotted the Fourier transformation of the charge current along the y-direction (the other two components are zero) as well as the magnetization of a typical Cr atom along the y-direction. As can be seen from Fig. 4(a), (b), (d), and (e), the response of CrSBr to the external laser pulse excitation contains odd harmonics in the electric current and as a consequence of inversion symmetry the even harmonics are canceled out. In all cases, high harmonics up to 7th harmonics are pronounced. While the higher harmonics are reduced in high fluence excitation, one can observe that the third harmonics survives when the excitation frequency is BG. In Fig. 4(c), the magnetization only contains the even harmonics. This is a consequence of SU(2) symmetry of the spin configuration along this direction [42] and can be seen from the magnetization dynamics in real time in Fig. 1(a) and (e) as well as Fig. 2(a) and (e). Nonetheless, the magnetization dynamics along the other directions (x and z) contain all even and odd harmonics (not shown here) which is driven by SOC interaction. However, the external magnetic field results in the breaking of the time-reversal symmetry and therefore the magnetization contains odd harmonics as well in this case (Fig. 4(f)). It should be noted that in this case, the z components of the electric current becomes nonzero due to the inverse spin Hall effect. In other word, the spin currents driven by the laser pulse, in the presence of the static external magnetic field, generate a perpendicular charge current along the z-direction.
Conclusion.—In this study, the ultrafast dynamics of semiconducting antiferromagnetic CrSBr is investigated using rt-TDDFT formalism. The ground state of the system was excited using a fs laser pulse with low and high fluence and below and above-gap central laser frequencies. We found high fluence excitation results in considerable demagnetization of the local magnetic moment of Cr atoms and it is almost independent of frequency. On the other hand, the dynamics of the magnetic moments of the Cr atoms strongly depends on the frequency in the low fluence excitation resulting in a slight demagnetization of Cr atoms with AG excitation. On the other hand, BG excitation enhances the magnetic moment of the Cr atoms; a peculiar observation which is just recently reported and is not yet carefully studied. In addition, an external applied magnetic field during the optical excitation triggers magnon generation in the nonequilibrium system. The microscopic mechanism of the demagnetization are found to be mostly intralayer and interlayer OISTR and SOC plays minor role. We observe hole generation in low fluence and electron generation in the high fluence excitation with both AG and BG frequencies. In addition, we observe odd and even harmonics generations in the charge current and magnetization, respectively.
Note added— Upon completing this work, we noticed a similar study the monolayer CrSBr [43].
Acknowledgements.
This research was supported by the US National Science Foundation (NSF) through the University of Delaware Materials Research Science and Engineering Center, DMR-2011824. The supercomputing time was provided by DARWIN (Delaware Advanced Research Workforce and Innovation Network), which is supported by NSF Grant No. MRI-1919839.References
- Du et al. [2016] K.-z. Du, X.-z. Wang, Y. Liu, P. Hu, M. I. B. Utama, C. K. Gan, Q. Xiong, and C. Kloc, Weak van der waals stacking, wide-range band gap, and raman study on ultrathin layers of metal phosphorus trichalcogenides, ACS nano 10, 1738 (2016).
- Kuo et al. [2016] C.-T. Kuo, M. Neumann, K. Balamurugan, H. J. Park, S. Kang, H. W. Shiu, J. H. Kang, B. H. Hong, M. Han, T. W. Noh, et al., Exfoliation and raman spectroscopic fingerprint of few-layer nips3 van der waals crystals, Scientific reports 6, 20904 (2016).
- Lin et al. [2016] M.-W. Lin, H. L. Zhuang, J. Yan, T. Z. Ward, A. A. Puretzky, C. M. Rouleau, Z. Gai, L. Liang, V. Meunier, B. G. Sumpter, et al., Ultrathin nanosheets of crsite 3: a semiconducting two-dimensional ferromagnetic material, Journal of Materials Chemistry C 4, 315 (2016).
- Lee et al. [2016] S. Lee, K.-Y. Choi, S. Lee, B. H. Park, and J.-G. Park, Tunneling transport of mono-and few-layers magnetic van der waals mnps3, APL Materials 4, 1 (2016).
- Huang et al. [2017] B. Huang, G. Clark, E. Navarro-Moratalla, D. R. Klein, R. Cheng, K. L. Seyler, D. Zhong, E. Schmidgall, M. A. McGuire, D. H. Cobden, et al., Layer-dependent ferromagnetism in a van der waals crystal down to the monolayer limit, Nature 546, 270 (2017).
- Gong et al. [2017] C. Gong, L. Li, Z. Li, H. Ji, A. Stern, Y. Xia, T. Cao, W. Bao, C. Wang, Y. Wang, et al., Discovery of intrinsic ferromagnetism in two-dimensional van der waals crystals, Nature 546, 265 (2017).
- Wu et al. [2019] M. Wu, Z. Li, T. Cao, and S. G. Louie, Physical origin of giant excitonic and magneto-optical responses in two-dimensional ferromagnetic insulators, Nature communications 10, 2371 (2019).
- Brennan et al. [2024] N. J. Brennan, C. A. Noble, J. Tang, M. E. Ziebel, and Y. J. Bae, Important elements of spin-exciton and magnon-exciton coupling, ACS Physical Chemistry Au 4, 322 (2024).
- Ahn et al. [2024] Y. Ahn, X. Guo, S. Son, Z. Sun, and L. Zhao, Progress and prospects in two-dimensional magnetism of van der waals materials, Progress in Quantum Electronics 93, 100498 (2024).
- Song and Xu [2024] T. Song and X. Xu, The future of 2d spintronics, Nature Reviews Electrical Engineering 1, 696 (2024).
- Geim and Grigorieva [2013] A. K. Geim and I. V. Grigorieva, Van der waals heterostructures, Nature 499, 419 (2013).
- Huang et al. [2018] B. Huang, G. Clark, D. R. Klein, D. MacNeill, E. Navarro-Moratalla, K. L. Seyler, N. Wilson, M. A. McGuire, D. H. Cobden, D. Xiao, et al., Electrical control of 2d magnetism in bilayer cri3, Nature nanotechnology 13, 544 (2018).
- Jiang et al. [2018] S. Jiang, J. Shan, and K. F. Mak, Electric-field switching of two-dimensional van der waals magnets, Nature materials 17, 406 (2018).
- Zhang et al. [2020] X.-X. Zhang, L. Li, D. Weber, J. Goldberger, K. F. Mak, and J. Shan, Gate-tunable spin waves in antiferromagnetic atomic bilayers, Nature materials 19, 838 (2020).
- Bae et al. [2022] Y. J. Bae, J. Wang, A. Scheie, J. Xu, D. G. Chica, G. M. Diederich, J. Cenker, M. E. Ziebel, Y. Bai, H. Ren, et al., Exciton-coupled coherent magnons in a 2d semiconductor, Nature 609, 282 (2022).
- Diederich et al. [2023] G. M. Diederich, J. Cenker, Y. Ren, J. Fonseca, D. G. Chica, Y. J. Bae, X. Zhu, X. Roy, T. Cao, D. Xiao, et al., Tunable interaction between excitons and hybridized magnons in a layered semiconductor, Nature Nanotechnology 18, 23 (2023).
- Zhang et al. [2022] P. Zhang, T.-F. Chung, Q. Li, S. Wang, Q. Wang, W. L. Huey, S. Yang, J. E. Goldberger, J. Yao, and X. Zhang, All-optical switching of magnetization in atomically thin cri3, Nature materials 21, 1373 (2022).
- Hu et al. [2022] H. Hu, N. Chen, H. Teng, R. Yu, Y. Qu, J. Sun, M. Xue, D. Hu, B. Wu, C. Li, et al., Doping-driven topological polaritons in graphene/-moo3 heterostructures, Nature Nanotechnology 17, 940 (2022).
- Ni et al. [2021] Z. Ni, A. Haglund, H. Wang, B. Xu, C. Bernhard, D. Mandrus, X. Qian, E. Mele, C. Kane, and L. Wu, Imaging the néel vector switching in the monolayer antiferromagnet mnpse3 with strain-controlled ising order, Nature nanotechnology 16, 782 (2021).
- Kang et al. [2020] S. Kang, K. Kim, B. H. Kim, J. Kim, K. I. Sim, J.-U. Lee, S. Lee, K. Park, S. Yun, T. Kim, et al., Coherent many-body exciton in van der waals antiferromagnet nips3, Nature 583, 785 (2020).
- Klaproth et al. [2023] T. Klaproth, S. Aswartham, Y. Shemerliuk, S. Selter, O. Janson, J. van den Brink, B. Büchner, M. Knupfer, S. Pazek, D. Mikhailova, et al., Origin of the magnetic exciton in the van der waals antiferromagnet nips3, Physical Review Letters 131, 256504 (2023).
- Ziebel et al. [2024] M. E. Ziebel, M. L. Feuer, J. Cox, X. Zhu, C. R. Dean, and X. Roy, Crsbr: an air-stable, two-dimensional magnetic semiconductor, Nano Letters 24, 4319 (2024).
- Lauk et al. [2020] N. Lauk, N. Sinclair, S. Barzanjeh, J. P. Covey, M. Saffman, M. Spiropulu, and C. Simon, Perspectives on quantum transduction, Quantum Science and Technology 5, 020501 (2020).
- Liu et al. [2023] S. Liu, I. A. Malik, V. L. Zhang, and T. Yu, Lightning the spin: Harnessing the potential of 2d magnets in opto-spintronics, Advanced Materials , 2306920 (2023).
- Gish et al. [2024] J. T. Gish, D. Lebedev, T. W. Song, V. K. Sangwan, and M. C. Hersam, Van der waals opto-spintronics, Nature Electronics 7, 336 (2024).
- Kirilyuk et al. [2010] A. Kirilyuk, A. V. Kimel, and T. Rasing, Ultrafast optical manipulation of magnetic order, Reviews of Modern Physics 82, 2731 (2010).
- Basov et al. [2017] D. Basov, R. Averitt, and D. Hsieh, Towards properties on demand in quantum materials, Nature materials 16, 1077 (2017).
- De La Torre et al. [2021] A. De La Torre, D. M. Kennes, M. Claassen, S. Gerber, J. W. McIver, and M. A. Sentef, Colloquium: Nonthermal pathways to ultrafast control in quantum materials, Reviews of Modern Physics 93, 041002 (2021).
- Beaurepaire et al. [1966] E. Beaurepaire, J.-C. Merle, A. Daunois, and J.-Y. Bigot, Ultrafast spin dynamics in ferromagnetic nickel, Physical review letters 76, 4250 (1966).
- Dirnberger et al. [2023] F. Dirnberger, J. Quan, R. Bushati, G. M. Diederich, M. Florian, J. Klein, K. Mosina, Z. Sofer, X. Xu, A. Kamra, et al., Magneto-optics in a van der waals magnet tuned by self-hybridized polaritons, Nature 620, 533 (2023).
- Sun et al. [2024] Y. Sun, F. Meng, C. Lee, A. Soll, H. Zhang, R. Ramesh, J. Yao, Z. Sofer, and J. Orenstein, Dipolar spin wave packet transport in a van der waals antiferromagnet, Nature Physics 20, 794–800 (2024).
- Diederich et al. [2025] G. M. Diederich, M. Nguyen, J. Cenker, J. Fonseca, S. Pumulo, Y. J. Bae, D. G. Chica, X. Roy, X. Zhu, D. Xiao, et al., Exciton dressing by extreme nonlinear magnons in a layered semiconductor, Nature Nanotechnology 20, 1 (2025).
- Huo et al. [2025] Y. Huo, S. Li, L. Yan, N. Li, S. Tretiak, and L. Zhou, Ultrafast magneto-lattice dynamics in two-dimensional crsbr driven by terahertz excitation, arXiv:2506.05058 (2025).
- Yang et al. [2021] K. Yang, G. Wang, L. Liu, D. Lu, and H. Wu, Triaxial magnetic anisotropy in the two-dimensional ferromagnetic semiconductor crsbr, Physical Review B 104, 144416 (2021).
- Dewhurst et al. [2016] J. Dewhurst, K. Krieger, S. Sharma, and E. Gross, An efficient algorithm for time propagation as applied to linearized augmented plane wave method, Comput. Phys. Commun. 209, 92 (2016).
- [36] The Elk Code, http://elk.sourceforge.net/.
- Klein et al. [2023] J. Klein, B. Pingault, M. Florian, M.-C. Heißenbüttel, A. Steinhoff, Z. Song, K. Torres, F. Dirnberger, J. B. Curtis, M. Weile, et al., The bulk van der waals layered magnet crsbr is a quasi-1d material, ACS nano 17, 5316 (2023).
- López-Paz et al. [2022] S. A. López-Paz, Z. Guguchia, V. Y. Pomjakushin, C. Witteveen, A. Cervellino, H. Luetkens, N. Casati, A. F. Morpurgo, and F. O. von Rohr, Dynamic magnetic crossover at the origin of the hidden-order in van der waals antiferromagnet crsbr, Nature Communications 13, 4745 (2022).
- Elliott et al. [2020] P. Elliott, N. Singh, K. Krieger, E. Gross, S. Sharma, and J. Dewhurst, The microscopic origin of spin-orbit mediated spin-flips, Journal of Magnetism and Magnetic Materials 502, 166473 (2020).
- Dewhurst et al. [2018] J. K. Dewhurst, P. Elliott, S. Shallcross, E. K. U. Gross, and S. Sharma, Laser-induced intersite spin transfer, Nano Lett. 18, 1842 (2018).
- Thorpe et al. [2023] A. Thorpe, N. Boroumand, A. Parks, E. Goulielmakis, and T. Brabec, High harmonic generation in solids: Real versus virtual transition channels, Physical Review B 107, 075135 (2023).
- Zhang et al. [2018] G. Zhang, M. Si, M. Murakami, Y. Bai, and T. F. George, Generating high-order optical and spin harmonics from ferromagnetic monolayers, Nature Communications 9, 3031 (2018).
- Sharma et al. [2025] S. Sharma, D. Gill, J. Krishna, E. Harris-Lee, J. K. Dewhurst, and S. Shallcross, Giant moment increase by ultrafast laser light, arXiv:2503.02396 (2025).